Two projectiles are fired from the ground level at equal speeds but different angles at an angle of 30 degree and the other one fired at 60° the projectile to hit the ground first would be the one fired at which degree neglect air resistance


1
Expert's answer
2020-10-25T18:27:38-0400

Equation of motion of projectile, launched at an angle "\theta" with initial speed "v_0" is "y(t) = v_0 \sin\theta t - \frac{g t^2}{2}". When it reaches the ground, "y(t) = 0", from where "t(v_0 \sin \theta -\frac{g t}{2}) = 0", so it takes "t = \frac{2 v_0 \sin \theta}{g}" to reach the ground. Since "\sin 30 = 0.5" and "\sin 60 \approx 0.87", and "v_0" is the same for both cases, projectile fired at an angle "30^{\circ}" will reach the ground faster.